OpenStudy (dmezzullo):

Use the distributive property to simplify the expression below: 5t − 6(1 + t) −t − 6 −t + 6 t + 6 6t − 6

OpenStudy (ajprincess):

Distributive property is a(b+c)=a*b+a*c. Using that u will get 5t − 6(1 + t) =5t-6*1-6*t =5t-6-6t can u simplify it nw @dmezzullo?

OpenStudy (dmezzullo):

no

OpenStudy (anonymous):

-t - 6